OpenStudy (anonymous):

Solve -2x2 +3x - 9 = 0

OpenStudy (anonymous):

hey @sam15389

OpenStudy (anonymous):

There are more than just one solution for this problem would you like both of them?

OpenStudy (raffle_snaffle):

do you have a ti-89

OpenStudy (raffle_snaffle):

do you have to justify your answer?

OpenStudy (anonymous):

First you have to simplify: x2+3x - 9

OpenStudy (raffle_snaffle):

I fanned you because you don't give the answer away.=)

OpenStudy (anonymous):

That's one thing I try my best never to do :D

OpenStudy (raffle_snaffle):

Sometimes they trick you! Lol

OpenStudy (anonymous):

Then you have to factor it out: \[x2+3x-9\] The first term is, x2 its coefficient is 1, the middle term is, +3x its coefficient is 3, the last term, "the constant", is -9 Step-1 : Multiply the coefficient of the first term by the constant 1 • -9 = -9 Step-2 : Find two factors of -9 whose sum equals the coefficient of the middle term, which is 3 . -9 + 1 = -8 -3 + 3 = 0 -1 + 9 = 8

OpenStudy (anonymous):

-2x2 +3x - 9 = 0 -4-9=-3x -13=-3x x=13/3 When you have a variable you don't want to deal with right a way just set it aside as I did. Then deal with what you can in front of you. After you've simplified your left side clean up both sides to solve for what you desired.

OpenStudy (anonymous):

depending on where you end up and how you perform through the succeeding steps then you will end up with one of two or both solutions

OpenStudy (radar):

Use the quadratic formula. The two roots are complex containing both real and imaginary values.

OpenStudy (campbell_st):

why not just use the general quadratic formula given the quadratic has no real roots. The discriminant < 0

OpenStudy (jhannybean):

Here's a tricky method to solve it, but it's still fun nonetheless. \[-2x^2 +3x -9=0\]Multiply your entire equation by -1\[2x^2 -3x+9=0\]Divide entire equation by 2 to get it in it's simplest form\[x^2 -\frac{3}{2}x +\frac{9}{2}=0\]Now we complete the square. \[\left( x^2 -\frac{3}{2}x\right)+\frac{9}{2} =0\]\[\left(x^2 -\frac{3}{2}x +\frac{9}{16}\right) +\frac{9}{2} -\frac{9}{16}=0\]\[\left( x-\frac{3}{4}\right)^2 +\frac{63}{16}=0\]\[\left(x-\frac{3}{4}\right)^2=-\frac{63}{16}\]\[x-\frac{3}{4} =\pm \sqrt{-\frac{63}{16}}\]\[x= \frac{3}{4} \pm \sqrt{-\frac{63}{16}}\]

OpenStudy (jhannybean):

\[x = \frac{3}{4} +i\frac{3\sqrt{7}}{4} \ ,\ \frac{3}{4}-i\frac{3\sqrt{7}}{4}\]

OpenStudy (radar):

Yes, completing the square is a practical way to solve it. In fact completing the square is one method to develop the "quadratic formula"